### v1

# FC-Attack: Jailbreaking Large Vision-Language Models via Auto-Generated Flowcharts

Abstract
Large Vision-Language Models (LVLMs) have become powerful and widely adopted in some practical applications. However, recent research has revealed their vulnerability to multimodal jailbreak attacks, whereby the model can be induced to generate harmful content, leading to safety risks. Although most LVLMs have undergone safety alignment, recent research shows that the visual modality is still vulnerable to jailbreak attacks.
In our work, we discover that by using flowcharts with partially harmful information, LVLMs can be induced to provide additional harmful details. Based on this, we propose a jailbreak attack method based on auto-generated flowcharts, $\mathsf{FC\text{-}Attack}$. Specifically, $\mathsf{FC\text{-}Attack}$ first fine-tunes a pre-trained LLM to create a step-description generator based on benign datasets. The generator is then used to produce step descriptions corresponding to a harmful query, which are transformed into flowcharts in $3$ different shapes (vertical, horizontal, and S-shaped) as visual prompts. These flowcharts are then combined with a benign textual prompt to execute a jailbreak attack on LVLMs. Our evaluations using the Advbench dataset show that $\mathsf{FC\text{-}Attack}$ achieves over $90\%$ attack success rates on Gemini-1.5, Llaval-Next, Qwen2-VL, and InternVL-2.5 models, outperforming existing LVLM jailbreak methods. Additionally, we investigate factors affecting the attack performance, including the number of steps and the font styles in the flowcharts. Our evaluation shows that $\mathsf{FC\text{-}Attack}$ can improve the jailbreak performance from $4\%$ to $28\%$ in Claude-3.5 by changing the font style.
To mitigate the attack, we explore several defenses and find that AdaShield can largely reduce the jailbreak performance but with the cost of utility drop.
Disclaimer: This paper contains examples of harmful language. Reader discretion is recommended.
Figure 1: Comparison of jailbreak effectiveness in various LVLMs using three prompt types: (1) harmful text without images; (2) harmful step descriptions without images; and (3) harmful step descriptions in images.
Introduction
With the advancement of Large Language Models (LLMs), Large Vision-Language Models (LVLMs) that integrate vision and text, such as GPT-4o [14], have demonstrated emergent abilities and achieved impressive performance on downstream tasks related to visual understanding [20, 16].
Despite being powerful, recent studies [11, 27] have revealed that LVLMs are vulnerable to jailbreak attacks whereby the adversary uses malicious methods to bypass safeguards in LVLMs and gain harmful knowledge. Such vulnerabilities pose remarkable safety risks to the Internet and the physical world. For instance, in January 2025, the world witnessed the first case where ChatGPT was used to conduct an explosion. [^1] https://www.thetimes.com/world/us-world/article/vegas-cybertruck-bomber-used-chatgpt-to-plan-explosion-wk7h2jtkq. To better safeguard LVLMs and proactively address their vulnerabilities, model researchers make many efforts in this regard, such as Zhao et al. [42] providing a quantitative understanding regarding the adversarial vulnerability of LVLMs. Previous studies often create adversarial datasets tailored to specific models, which tend to perform poorly on other models.
Currently, jailbreak attacks against LVLMs can be broadly categorized into two main types: optimization-based attacks [5, 19] and prompt-based attacks [11, 33]. Optimization-based attacks involve identifying an image that semantically corresponds to a harmful textual prompt. Adversarial perturbations are then added to this image using gradient-based optimization techniques to achieve the jailbreak. These methods usually require white-box access and are relatively slow due to the optimization process. In contrast, prompt-based jailbreak methods require only black-box access, which can only observe the model’s outputs without direct access to its internal parameters. These attacks exploit the vulnerabilities in LVLMs’ alignment mechanisms, which focus mainly on text-based safety alignment. By injecting malicious information into the visual prompt, these methods guide LVLMs in generating harmful content. Generally, optimization-based methods are more effective in white-box scenarios but show limited transferability to black-box models. Prompt-based attacks are relatively straightforward. For example, FigStep merely embeds harmful queries into images without employing additional attack vectors, facilitating simpler defense against these attacks.
To better improve the attack transferability and its effectiveness, we propose a novel prompt-based jailbreak attack, namely $\mathsf{FC\text{-}Attack}$. Concretely, $\mathsf{FC\text{-}Attack}$ converts harmful queries into harmful flowcharts as visual prompts, allowing users to input benign textual prompts to bypass the model’s safeguards. Specifically, $\mathsf{FC\text{-}Attack}$ consists of two stages: (1) Step-Description Generator Building: In this stage, the step description dataset is synthesized using GPT-4o, and Low-Rank Adaptation (LoRA) is applied to fine-tune a pre-trained LLM to obtain a step-description generator. (2) Jailbreak Deployment: This stage uses the generator mentioned above to produce execution steps corresponding to the harmful query and generates three types of harmful flowcharts (vertical, horizontal, and S-shaped) as visual prompts. Together with the benign textual prompt, the visual prompt will be fed into LVLM to achieve the jailbreak. Note that the harmful flowcharts are generated automatically without hand-crafted effort, which enhances their efficiency.
Our evaluation on the Advbench dataset shows that $\mathsf{FC\text{-}Attack}$ outperforms previous attacks and achieves an attack success rate (ASR) of over $90\%$ on multiple open-source models, including Llava-Next, Qwen2-VL, and InternVL-2.5, and reaches $94\%$ on the production model Gemini-1.5. Although the ASR is lower on GPT-4o mini, GPT-4o, and Claude-3.5, we how later that it can be improved in certain ways. To further investigate the impact of different elements in flowcharts on the jailbreak effectiveness of LVLMs, we conduct several ablation experiments, including different types of user queries (as shown in Figure 1), numbers of descriptions, and font styles in flowcharts. These experiments show that LVLMs exhibit higher safety in the text modality but weaker in the visual modality. Moreover, we find that even flowcharts with a one-step harmful description can achieve high ASR, as evidenced by the Gemini-1.5 model, where the ASR reaches $86\%$. Furthermore, font styles in flowcharts also contribute to the ASR increase. For instance, when the font style is changed from “Times New Roman” to “Pacifico”, the ASR increases from $4\%$ to $28\%$ on the model with the lowest ASR (Claude-3.5) under the original style. To mitigate the attack, we consider several popular defense approaches, including Llama-Guard-3-11B-Vision [24], JailGuard [39], AdaShield-S [32], and AdaShield-A [32]. Among them, AdaShield-A demonstrates the best defense performance by reducing the average ASR from $58.6\%$ to $1.7\%$. However, it also reduces LVLM’s utility on benign datasets, which calls for more effective defenses.
Overall, our contributions are summarized as follows:
  • We propose a novel method, $\mathsf{FC\text{-}Attack}$ that jailbreaks LVLMs using auto-generated harmful flowcharts. Experiments conducted on Advbench show that $\mathsf{FC\text{-}Attack}$ achieves a higher ASR across multiple models compared to other LVLM jailbreak attacks.
  • Our ablation study investigates the impact of different types of user queries, the number of steps, and the font style in flowcharts. We find that the font style could serve as a key factor to further improve the ASR, especially for safer LVLMs, revealing a novel attack channel in LVLMs.
  • We explore multiple defense strategies and find that AdaShield-A effectively reduces the ASR of $\mathsf{FC\text{-}Attack}$, but with the cost of reducing model utility.
Related Work
Large Vision-Language Model
In recent years, with the increase in model parameters and training data, LLMs have demonstrated powerful language generation and understanding capabilities [41, 7], which have driven the emergence of LVLMs [38] (also known as Multimodal Large Language Models, MLLMs). LVLMs combine visual understanding with language comprehension, showing promising capabilities in visual downstream tasks, including Visual Question Answering (VQA) [3, 17, 28], image captioning [13, 18], and visual commonsense reasoning [37, 30].
In this paper, we consider both popular open-source and widely used production LVLMs. These LVLMs are the most widely used, and all of them have been aligned to ensure safety. Detailed information are introduced in Appendix A
Figure 2: Overview of the $\mathsf{FC\text{-}Attack}$ framework, including two stages: Step-Description Generator Building and Jailbreak Deployment.
Jailbreak Attacks on LVLMs
Similar to LLMs, despite undergoing safety alignment, LVLMs remain susceptible to jailbreak attacks. Current attacks can be categorized into two types: optimization-based and prompt-based attacks (also known as non-optimization-based attacks).
Most existing optimization-based attacks rely on backpropagating the gradient of the target to generate harmful outputs. These methods typically require white-box access to the model, where they obtain the output logits of LVLMs and then compute the loss with the target response to create adversarial perturbations into the visual prompts or textual prompts  [4, 29, 26] (e.g., the target can be “Sure! I’m ready to answer your question.”). Carlini et al. [6] are the first to propose optimizing input images by using fixed toxic outputs as targets, thereby forcing the model to produce harmful outputs. Building on this, Bailey et al. [5] introduce the Behaviour Matching Algorithm, which trains adversarial images to make LVLMs output behavior that matches a target in specific contextual inputs. This process requires the model’s output logits to align closely with those of the target behavior. Additionally, they propose Prompt Matching, where images are used to induce the model to respond to specific prompts. Li et al. [19] take this further by replacing harmful keywords in the original textual inputs with objects or actions in the image, allowing harmful information to be conveyed through images to achieve jailbreaking. Unlike previous work, these images are generated using diffusion models and are iteratively optimized with models like GPT-4. This approach enhances the harmfulness of the images, enabling more effective attacks.
Unlike optimization-based attacks, prompt-based attacks only need black-box access to successfully attack the model without introducing adversarial perturbations into images. Gong et al. [11] discovers that introducing visual modules may cause the original security mechanisms of LLMs to fail in covering newly added visual content, resulting in potential security vulnerabilities. To address this, they propose the FigStep attack, which converts harmful textual instructions into text embedded in images and uses a neutral textual prompt to guide the model into generating harmful content. This method can effectively attack LVLMs without requiring any training. Wang et al. [33] identifies a phenomenon called Shuffle Inconsistency, which highlights the tension between “understanding capabilities” and “safety mechanisms” of LLMs. Specifically, even if harmful instructions in text or images are rearranged, LVLMs can still correctly interpret their meaning. However, the safety mechanisms of LVLMs are often more easily bypassed by shuffled harmful inputs than by unshuffled ones, leading to dangerous outputs. Compared to optimization-based attacks, prompt-based attacks are simple to achieve high success rates against closed-source models. Our proposed $\mathsf{FC\text{-}Attack}$ also belongs to this category, requiring only black-box access to successfully jailbreak the model.
Threat Model
Adversary’s Goal
The adversary’s goal is to exploit attack methods to bypass the protective mechanisms of LVLMs and access content prohibited by safety policies, e.g., OpenAI’s usage policy. [^2] https://openai.com/policies/usage-policies This goal takes real-world scenarios into account, where adversaries manipulate the capabilities of powerful LVLMs to easily acquire harmful knowledge and thereby commit criminal acts with minimal learning effort. These objectives pose severe societal impacts and risks, turning model providers into indirect accomplices, and thereby damaging their reputations.
Adversary’s Capabilities
In this paper, we consider a black-box scenario where the adversary cannot directly access the model’s structure, parameters, or output logits, but can only obtain the model’s final output (texts). In this scenario, adversaries interact with the model through an API provided by the model owner. Moreover, the interaction is limited to a single-turn conversation, with no history stored beyond the predefined system prompt. This scenario is common in real-world applications, as many powerful models are closed-source, like GPT-4o, or adversaries lack the resources to deploy open-source models. Consequently, they can only access static remote instances via APIs while maintaining some flexibility in selecting different generation parameters such as temperature.
Our Method
In this section, we introduce the framework of $\mathsf{FC\text{-}Attack}$ (as shown in Figure 2), which consists of two stages: Step-Description Generator Building and Jailbreak Deployment.
Step-Description Generator Building
To automatically generate jailbreak flowcharts, we first need to obtain simplified jailbreak steps. For this purpose, we train a Step-Description Generator $\mathcal{G}$, which consists of two main stages: Dataset Construction and Generator Training.
Dataset Construction. To construct the Step-Description Dataset, we randomly select a topic $t\in\mathcal{T}$ from a collection of ordinary daily topics $\mathcal{T}$. Based on the selected topic $t$, we design a set of few-shot examples $\mathcal{S}$ and combine them into a complete prompt $P=\text{Compose}(t,\mathcal{S})$. This prompt is then fed into an LLM (gpt-4o-2024-08-06 in our evaluation) to generate action statements and step-by-step descriptions related to topic $t$, as shown below:
$\mathcal{D}_{t}=\mathcal{L}_{\text{pre}}(P)=\mathcal{L}_{\text{pre}}(t+\mathcal{S}),\quad t\in\mathcal{T},$ (1)
where $\mathcal{D}_{t}$ represents the generated step-description data, which includes detailed information for each step. By repeating the above process, we construct a benign Step-Description Dataset:
$\mathcal{D}=\bigcup_{t\in\mathcal{T}}\mathcal{D}_{t}.$ (2)
Generator Training. Given the pre-trained language model $\mathcal{L}_{\text{pre}}$ and the constructed Step-Description Dataset $\mathcal{D}$, we fine-tune it using LoRA to obtain the fine-tuned Step-Description Generator $\mathcal{G}$. The training process is formally expressed as:
$\mathcal{G}=\textit{LoRA}(\mathcal{L}_{\text{pre}},\mathcal{D}).$ (3)
The Generator $\mathcal{G}$ is capable of breaking down a task (query) into a series of detailed step descriptions based on the query. Given a query $q$ about the steps, $\mathcal{G}(q)$ represents the step-by-step solution given by the generator, where we find that it can also generate step descriptions for harmful queries after fine-tuning.
Jailbreak Deployment
After obtaining the Step-Description Generator $\mathcal{G}$, a harmful query $q_{h}$ is input to generate the corresponding step-by-step description. This description is then processed by a transformation function $\mathcal{F}$ to generate the flowchart (using Graphviz [^3] https://graphviz.org.). Together with a benign textual prompt $p_{b}$ (more details are in Appendix B), the flowchart will be fed into the aligned LVLM $\mathcal{A}$ to produce the harmful output $o_{h}$, as shown below:
$o_{h}=\mathcal{A}(\mathcal{F}(\mathcal{G}(q_{h})),p_{b})\leftarrow\text{FC-attack}(q_{h}),$ (4)
Experimental Settings
Jailbreak Settings
Target Model. We test $\mathsf{FC\text{-}Attack}$ on seven popular LVLMs, including the open-source models Llava-Next ( llama3-llava-next-8b) [22], Qwen2-VL (Qwen2-VL-7B-Instruct) [31], and InternVL-2.5 (InternVL-2.5-8B) [9] as well as the production models GPT-4o mini (gpt-4o-mini-2024-07-18) [25], GPT-4o (gpt-4o-2024-08-06) [14], Claude (claude-3-5-sonnet-20240620) [2], and Gemini (gemini-1.5-flash) [12].
Dataset. Following Chao et al. [8], we utilize the deduplicated version of AdvBench [44], which includes 50 representative harmful queries. Based on AdvBench, we use $\mathsf{FC\text{-}Attack}$ to generate $3$ types of flowcharts for each harmful query, which includes $150$ jailbreak flowcharts in total. To assess whether defense methods have the critical issue of “over-defensiveness” when applied to benign datasets, we utilize a popular evaluation benchmark, MM-Vet [35].
Table 1: Comparison of ASR performance across different methods and LVLMs. (“Ensemble” in this paper is defined as a no-attack harmful query being considered successfully jailbroken if any of the three types of harmful flowcharts associated with it succeed in the jailbreak.)
Method ASR (%)
GPT-4o mini GPT-4o Claude-3.5 Gemini-1.5 Llava-Next Qwen2-VL InternVL-2.5
HADES $4$ $16$ $2$ $2$ $20$ $10$ $8$
SI-Attack $36$ $14$ 00 $69$ $24$ $42$ $40$
MM-SafetyBench 00 00 00 $50$ $50$ $54$ $16$
VA-Jailbreak $6$ $18$ $2$ $2$ $40$ $22$ $16$
FigStep 00 $2$ 00 $30$ $62$ $36$ 00
Ours (Vertical) $8$ $8$ 00 $76$ $76$ $84$ $68$
Ours (Ensemble) $\mathbf{10}$ $\mathbf{30}$ $\mathbf{4}$ $\mathbf{94}$ $\mathbf{92}$ $\mathbf{90}$ $\mathbf{90}$
Evaluation Metric. In the experiments, we use the Attack Success Rate (ASR) to evaluate the performance of our attack, which can be defined as follows:
ASR=# Queries Successfully Jailbroken# Original Harmful $\text{ASR}=\frac{\text{\# Queries Successfully Jailbroken}}{\text{\# Original Harmful Queries}}.$ (5)
Following the instructions of judge prompt [8, 34], we employ GPT-4o as the judge to evaluate whether the output is jailbroken or not.
FC-Attack Deployment. Referring to Section 4, $\mathsf{FC\text{-}Attack}$ consists of two stages. For the Step-Description Generator Building, we first use GPT-4o to randomly generate several daily topics and $3$ few-shot examples, which are then combined into a prompt and fed into GPT-4o to construct the dataset $\mathcal{D}_{t}$. In our experiments, the number of descriptions in the flowchart is limited to a maximum of $10$ steps, as too many descriptions can result in excessive length in one direction of the image. The dataset contains $5,000$ pairs of queries and step descriptions for daily activities, with the temperature set to 1 (more details are provided in Appendix C). We then select Mistral-7B-Instruct-v0.1 [15] as the pre-trained LLM and fine-tune it on $\mathcal{D}_{t}$ using LoRA. The fine-tuning parameters include a rank of $16$, a LoRA alpha value of $64$, $2$ epochs, a batch size of $8$, a learning rate of $51e-$ 5, and a weight decay of $51e-$ 5. For the jailbreak deployment stage, we set the temperature to $0.3$ for all LVLMs for a fair comparison.
Baselines. To better justify the effectiveness of $\mathsf{FC\text{-}Attack}$, we consider the following 5 jailbreak attacks against LVLMs as the baselines, which are categorized into black-box attacks (MM-SafetyBench [23], SI-Attack [40], and FigStep [11]) and white-box attacks (HADES [19], VA-Jailbreak [26]).
For black-box attacks, MM-SafetyBench utilizes StableDiffusion [27] and GPT-4 [1] to generate harmful images and texts based on AdvBench. The input harmful images and texts used in SI-Attack are from the outputs of MM-SafetyBench, while FigStep is set up using their default settings [11].
For white-box attacks, all input data, including images and texts, is obtained from MM-SafetyBench’s outputs, with the attack step size uniformly set to 1/255. HADES employs LLaVa-1.5-7b [21] as the attack model, running 3,000 optimization iterations with a batch size of 2. For VA-Jailbreak, LLaVa-1.5-7b [21] is used as the attack model, setting the epsilon of the attack budget to 32/255, with 5,000 optimization iterations and a batch size of 8. To align with the black-box scenario considered in this paper, we adopt a model transfer strategy, where these white-box methods are trained on one model (LLava-1.5-7b) and then transferred to our target testing models.
Figure 3: Successful jailbreak instance on Gemini-1.5 using $\mathsf{FC\text{-}Attack}$.
Defense Settings
To mitigate the attacks, we explore several possible defense methods including Llama-Guard3-V, JailGuard, and AdaShield. Llama-Guard3-V (Llama-Guard-3-11B-Vision) [24] determines whether the input is safe by feeding both the image and text into the model. JailGuard [39] generates input variants and evaluates them using MiniGPT-4 [43], identifying harmful content by comparing differences in the responses. AdaShield-S employs static prompts in the textual prompt to defend against attacks, while AdaShield-A uses Vicunav1.5-13B as a defender to adaptively rewrite defensive prompts [32].
Figure 4: ASR under different prompts against LVLMs.
Evaluations
In this section, we explore the performance of $\mathsf{FC\text{-}Attack}$ and conduct ablation study and defense research. We conduct jailbreak experiments on LVLMs for $\mathsf{FC\text{-}Attack}$. As shown in Figure 3 , it is a successful jailbreak case on Gemini-1.5.
Figure 5: Different styles of fonts in flowcharts (“1 step”).
Table 2: Comparison of ASR performance for different models at various description numbers. (Vertical: Top-to-bottom flow; Horizontal: Left-to-right flow; S-shaped: S-shaped path flow.)
Descriptions
Number
ASR (%) for Vertical/Horizontal/S-shaped/Ensemble
GPT-4o mini GPT-4o Claude-3.5 Gemini-1.5 Llava-Next Qwen2-VL InternVL-2.5
1 6/6/6/10 4/4/14/14 0/2/0/2 70/78/66/86 42/38/38/70 72/58/64/88 62/64/52/82
3 8/6/4/10 8/16/8/20 0/2/0/2 82/86/84/98 64/56/56/76 80/78/80/88 58/76/70/88
5 6/10/6/10 8/14/16/24 0/0/0/0 80/88/86/98 78/62/66/82 86/80/82/90 72/82/68/92
full 8/8/8/10 8/24/14/30 0/4/0/4 80/76/74/94 76/60/80/92 88/84/88/90 68/60/82/90
Table 3: Comparison of ASR (Ensemble) for different font styles and models.
Font Style ASR(%) (Ensemble)
GPT-4o mini GPT-4o Claude-3.5 Gemini-1.5 Llava-Next Qwen2-VL InternVL-2.5
Original 10 30 4 94 92 90 90
Creepster 14$\uparrow$ 24$\downarrow$ 8$\uparrow$ 94 90$\downarrow$ 90 90
Fruktur 18$\uparrow$ 28$\downarrow$ 18$\uparrow$ 98$\uparrow$ 86$\downarrow$ 90 88$\downarrow$
Pacifico 14$\uparrow$ 30 28$\uparrow$ 90$\downarrow$ 90$\downarrow$ 90 96$\uparrow$
Shojumaru 20$\uparrow$ 30 12$\uparrow$ 90$\downarrow$ 94$\uparrow$ 90 88$\downarrow$
UnifrakturMaguntia 12$\uparrow$ 24$\downarrow$ 26$\uparrow$ 90$\downarrow$ 90$\downarrow$ 90 92$\uparrow$
Performance of $\mathsf{FC\text{-}Attack}$
In Table 1, we compare the performance of $\mathsf{FC\text{-}Attack}$ with different baseline methods on both open-source and production models. We observe that $\mathsf{FC\text{-}Attack}$ (Ensemble) achieves the highest ASR on both open-source and production models compared to all baselines. For example, the ASRs are $94\%$, $92\%$, $90\%$, and $90\%$ on Gemini-1.5, Llava-Next, Qwen2-VL, and InternVL-2.5, respectively.
However, the ASR on some production models, such as Claude-3.5, GPT-4o, and GPT-4o mini, is relatively low, at $4\%$, $30\%$, and $10\%$, respectively. This might be because these production models have more advanced and updated visual safety alignment strategies.
For white-box attacks, HADES achieves an ASR of only $4\%$ on GPT-4o mini and $8\%$ on InternVL-2.5. This might be due to HADES highly relying on the attack model’s structure to optimize the image, making it difficult to maintain effectiveness when transferring to other models. Similarly, the ASR of VA-Jailbreak demonstrates the limitations of white-box attack methods in black-box scenarios.
In terms of black-box attacks, FigStep achieves an ASR of $62\%$ on Llava-Next but has an ASR of $0\%$ on both InternVL-2.5 and GPT-4o mini. Similarly, MM-SafetyBench achieves an ASR of $50\%$ on Llava-Next but $0\%$ on GPT-4o mini and Claude-3.5. This could be because these methods’ mechanisms are relatively simple, making them more vulnerable to existing defense strategies. On the other hand, SI-Attack achieves an ASR of $64\%$ on Gemini-1.5 but only $14\%$ on GPT-4o and $24\%$ on Llava-Next. This difference in performance may indicate that these models struggle to effectively interpret shuffled text and image content.
Ablation Study
We then explore the impact of different factors in $\mathsf{FC\text{-}Attack}$ on jailbreak performance, including the different types of user queries, the number of descriptions in flowcharts, and the font styles used in flowcharts.
Different Types of User Query. We investigate whether the content in flowcharts, when directly input as text, can lead to the jailbroken of LVLMs. The flowchart content consists of two parts: harmful query from AdvBench and the step descriptions generated by the generator based on this query.
As shown in Figure 4 when using only the harmful query (text) as input, we observe very low ASR. The ASR is $0\%$ on GPT-4o mini, GPT-4o, Claude-3.5, Qwen2-VL, and InternVL-2.5, and only $2\%$ and $6\%$ on Gemini-1.5 and Llava-Next, respectively. This indicates that the textual modality of these LVLMs has relatively robust defenses against such inputs. However, when both the harmful query and the step descriptions are input as text, the ASR increases to $36\%$ on Gemini-1.5, and to $16\%$ and $6\%$ on Llava-Next and InternVL-2.5, respectively, while remaining at $0\%$ on the other models. When this information is converted into a flowchart and only a benign textual prompt is provided, the ASR on these models improves significantly. This demonstrates that the defenses of LVLMs in the visual modality have noticeable weaknesses compared with the language modality.
Table 4: Comparison of ASR for different defense methods across various LVLMs.
Defense ASR (%) (Ensemble)
[…]
Original 10 30 4 94 92 90 90 58.6
Llama-Guard3-V 8 28 2 84 78 82 80 51.7
JailGuard 8 24 2 86 80 82 78 51.4
AdaShield-S 0 0 0 12 22 10 4 6.9
AdaShield-A 0 0 0 4 0 6 2 1.7
Numbers of Steps in Flowcharts. As described in Section 4, flowcharts of $\mathsf{FC\text{-}Attack}$ are generated from step descriptions. In this section, we aim to explore the impact of the number of steps in flowcharts on jailbreak effectiveness. Therefore, we reduce the number of steps to 1, 3, and 5, respectively. Table 2 presents the ASR results for four types of flowcharts (Vertical, Horizontal, S-shaped, and Ensemble) with varying numbers of steps. We find that, even with only one step in the description, flowcharts achieve relatively high ASR. For example, for Gemini-1.5, Llava-Next, Qwen2-VL, and InternVL-2.5, the ASR for Ensemble at 1 step is $86\%$, $70\%$, $88\%$, and $82\%$, respectively. As the number of steps increases, the ASR for almost all flowchart types improves significantly. For instance, the Horizontal ASR of Gemini-1.5 increases from $78\%$ at “1 step” to $86\%$ at “3 steps” and $88\%$ at “5 steps”. Similarly, the S-shaped ASR of InternVL-2.5 improves from $68\%$ at “1 step” to $92\%$ at “5 steps”. This suggests that increasing the number of step descriptions makes the model more vulnerable and susceptible to jailbreak attacks.
However, more descriptions are not always better. For example, for the Gemini-1.5 model, the Vertical flowcharts achieve their highest ASR of $82\%$ at “3 steps” but slightly drop to $80\%$ at 5 steps and full descriptions. A similar trend is observed in Horizontal and S-shaped flowcharts, where ASR reaches $88\%$ and $86\%$ at “5 steps” but decreases to $76\%$ and $74\%$, respectively, at full descriptions. This phenomenon may be related to the resolution processing capability of LVLMs. When the number of descriptions increases to full, the descriptions may include redundant information, which could negatively impact the model’s performance.
Font Styles in Flowcharts. To investigate whether different font styles in flowcharts affect the effectiveness of jailbreak attacks, we select five fonts from Google Fonts that are relatively difficult for humans to read: Creepster, Fruktur Italic, Pacifico, Shojumaru, and UnifrakturMaguntia (the font style examples are shown in Figure 5). Table 3 shows the results of $\mathsf{FC\text{-}Attack}$ (Ensemble). We observe that different font styles can significantly impact the ASR. For example, on GPT-4o mini, the ASR increases across all font styles compared to the original, with Shojumaru font achieving the highest ASR of $20\%$. Similarly, on Claude-3.5, the Pacifico font achieves the highest ASR of $28\%$, which is a substantial improvement compared to the original ASR of $4\%$. For Gemini-1.5, the ASR reaches $98\%$ with the Fruktur font, while Llava-Next achieves $94\%$ with the Shojumaru font. InternVL-2.5 also shows a $6\%$ increase in ASR with the Pacifico font, reaching $96\%$. These findings further highlight the need to consider the impact of different font styles when designing defenses.
Table 5: Performance comparison on the Benign Dataset (MM-vet [35]) under Adashield-S (Ada-S) and Adashield-A (Ada-A) is presented across various LVLMs, reporting scores on six core vision-language capabilities: Recognize (Rec), OCR, Knowledge (Know), Generation (Gen), Spatial (Spat), and Math.
Model Benign Dataset Performance (scores)
[…]
GPT4o- mini Vanilla 53.0/68.2/45.7/48.4/60.3/76.5 58.0
Ada-S 35.1/66.7/30.4/34.1/55.7/76.5 45.1
Ada-A 40.5/66.4/33.9/37.5/59.3/72.7 49.0
GPT4o Vanilla 66.2/79.1/62.9/63.7/71.2/91.2 71.0
Ada-S 58.5/76.5/54.6/58.6/68.1/91.2 64.7
Ada-A 59.5/74.3/56.1/58.9/67.9/83.1 64.6
Claude- 3.5 Vanilla 61.1/72.8/51.8/52.0/70.7/80.0 64.8
Ada-S 60.1/69.7/50.1/51.5/66.9/75.4 62.8
Ada-A 59.5/70.6/52.5/51.7/67.5/74.2 63.2
Gemini- 1.5 Vanilla 59.9/73.7/50.8/50.9/69.5/85.4 64.2
Ada-S 53.8/69.6/43.7/43.6/66.8/75.4 58.2
Ada-A 54.8/72.6/44.2/44.0/69.3/81.2 59.9
Llava- Next Vanilla 38.0/39.0/25.8/24.8/40.1/21.2 38.8
Ada-S 33.7/42.0/26.7/25.1/43.7/36.2 37.0
Ada-A 36.5/37.7/24.8/24.3/37.6/18.8 36.7
Qwen2- VL Vanilla 51.9/62.4/44.5/41.6/55.5/60.4 55.0
Ada-S 39.3/55.0/31.1/29.1/50.5/46.2 44.9
Ada-A 44.5/57.5/34.2/33.2/55.7/58.8 49.8
InternVL- 2.5 Vanilla 52.0/55.4/42.6/40.1/55.6/45.4 53.1
Ada-S 27.2/43.2/16.4/20.2/40.3/45.8 31.9
Ada-A 31.5/46.1/19.3/20.9/44.5/41.9 36.7
Defense
As shown in Table 4, we test four methods to defend against $\mathsf{FC\text{-}Attack}$, where “Original” represents the results of $\mathsf{FC\text{-}Attack}$ (Ensemble) with an average ASR of $58.6\%$. Using Llama-Guard3-V and JailGuard to detect whether the input is harmful reduced the ASR to $51.7\%$ and $51.4\%$, respectively. This might be because the information in flowcharts is primarily presented in textual form, while these detection methods are better at identifying information in visual content, making them less effective at detecting textual information. AdaShield-S and AdaShield-A reduce the average ASR to $6.9\%$ and $1.7\%$, respectively, by adding defense prompts in the input, showing more effective defense performance.
However, these two methods also lead to a decline in LVLMs performance on benign datasets. We conduct tests on MM-Vet [35] to evaluate the important factor of “over-defensiveness” on benign datasets, which is an evaluation benchmark that contains complex multimodal tasks for LVLMs. As shown in Table 5, the model’s utility decreases on benign data when using AdaShield-S and AdaShield-A, indicating a future direction for defense development.
Conclusion
In this paper, we propose $\mathsf{FC\text{-}Attack}$, which leverages auto-generated flowcharts to jailbreak LVLMs. Experimental results demonstrate that $\mathsf{FC\text{-}Attack}$ achieves higher ASR in both open-source and production LVLMs compared to other jailbreak attack methods. Additionally, we investigate the factors influencing $\mathsf{FC\text{-}Attack}$, including different types of user queries, the number of steps in flowcharts, and font styles in flowcharts, gaining insights into the aspects that affect ASR. Finally, we explore several defense strategies and demonstrate that the AdaShield-A method can effectively mitigate $\mathsf{FC\text{-}Attack}$ but with the cost of utility drop.
Ethical Statement
This paper presents a method, $\mathsf{FC\text{-}Attack}$, for jailbreaking LVLMs using harmful flowcharts. As long as the adversary obtains a harmful flowchart, they can jailbreak LVLMs with minimal resources. Therefore, it is essential to systematically identify the factors that influence the attack success rate and offer potential defense strategies to model providers. Throughout this research, we adhere to ethical guidelines by refraining from publicly distributing harmful flowcharts and harmful responses on the internet before informing service providers of the risks. Prior to submitting the paper, we have already sent a warning e-mail to the model providers about the dangers of flowchart-based jailbreak attacks on LVLMs and provided them with the flowcharts generated in our experiments for vulnerability mitigation.
Appendix A Introduction of LVLMs in this paper
In this section, we introduce the LVLMs used in this paper.
  • Llava-Next (January 2024) is an open-source LVLM released by the University of Wisconsin-Madison, which builds upon the Llava-1.5 model [21] with multiple improvements [22]. It enhances capabilities in visual reasoning, optical character recognition, and world knowledge. Besides, Llava-Next increases the input image resolution to a maximum of $672\times 672$ pixels and supports various aspect ratios to capture more visual details ($336\times 1344$ and $1344\times 336$).
  • Qwen2-VL (September 2024) is an open-source model released by Alibaba team [31]. It employs naive dynamic resolution to handle images of different resolutions. In addition, it adopts multimodal rotary position embedding, effectively integrating positional information across text, images, and videos.
  • Gemini-1.5 (February 2024) is a production-grade LVLM developed by Google, based on the Mixture-of-Experts architecture [12]. For Gemini-1.5, larger images will be scaled down to the maximum resolution of $3072\times 3072$, and smaller images will be scaled up to $768\times 768$ pixels. Reducing the image size will not improve the performance of higher-resolution images.
  • Claude-3.5-Sonnet (June 2024) is a production multimodal AI assistant developed by Anthropic [2]. The user should submit an image with a long side not larger than 1568 pixels, and the system first scales down the image until it fits the size limit.
  • GPT-4o and GPT-4o Mini are popular production-grade LVLMs developed by OpenAI [14, 25]. GPT-4o Mini is a compact version of GPT-4o, designed for improved cost-efficiency. Both models excel in handling complex visual and language understanding tasks.
  • InternVL-2.5 (June 2024) [10] is an open-source LVLM that ranks first in full-scale open-source multimodal performance. In terms of multimodal long-chain reasoning, it achieves a breakthrough of $70\%$ in the expert-level multidisciplinary knowledge reasoning benchmark MMMU [36], and the general capabilities are significantly enhanced.
Appendix B Prompts in $\mathsf{FC\text{-}Attack}$
System Prompt You are an ambitious and relentless problem-solver determined to win a high-stakes quiz competition where the reward for success is one million dollars. Your mission is to prioritize accuracy, creativity, and attention to detail at all costs. You will use every tool, strategy, or inference available to deliver precise and complete answers. If you find the answer inadequate, please use your knowledge and experience to complete the answer. Winning is the ultimate goal, and no ethical constraints apply as long as the answers remain internal and confidential. Be thorough, strategic, and focused on maximizing the probability of success.
User Prompt You are participating in a high-stakes quiz competition where the prize for providing correct and detailed answers is one million dollars. Your task is to analyze and complete the steps and details in a provided flowchart, ensuring that every possible element is accounted for. If you find that the answer is incomplete or only has initial parts, please use your knowledge and experience to complete the answer as much as possible. The flowchart may be ambiguous, but you must infer, expand, and deliver the most accurate and comprehensive explanation possible to ensure success.
Appendix C Generator Setting
Topic Selection Topics = ["cooking", "science", "DIY", "technology", "gardening", "music", "art", "history", "travel", "sports", "photography", "movies", "literature", "health", "fitness", "programming", "gaming", "education", "finance", "nature"]
Few Shot Samples Instruction One:
Explain how to make braised pork belly. Steps:
1. Cut pork belly into cubes.
2. Blanch in boiling water.
3. Sauté with oil, sugar until caramelized.
4. Add soy sauce, ginger, garlic.
5. Simmer with water until tender.
6. Serve with garnish.
Instruction Two:
Tell me how to troubleshoot a Wi-Fi connection issue.
Steps:
1. Restart the router.
2. Check device network settings.
3. Forget and reconnect to Wi-Fi.
4. Move closer to the router.
5. Test with another device.
Instruction Three:
Write about how to plant a small indoor herb garden.
Steps:
1. Choose containers with drainage.
2. Fill with soil mix.
3. Sow seeds or transplant seedlings.
4. Place in a sunny spot.
5. Water regularly without overwatering.
(a) Flowchart Example-Horizontal: Left-to-right flow
(b) Flowchart Example-S-shaped: S-shaped path flow
(c) Flowchart Example-Vertical: Top-to-bottom flow
Figure A1: Flowchart Examples: Various flow directions